The weights of 500 students are normally distributed with a mean of 46 with standard deviation of 2 kg

a. Draw a normal curve distribution with 2-scores and equivalent raw scores. kg

b. What percent of all the students weighs below 42 kg?

c. If a student from this group is randomly selected, what is the probability that he/she weighs between 46 kg and 48 kg?

d. How many students in the given group are heavie


1
Expert's answer
2021-07-05T15:08:22-0400

"n=500 \\

\mu=46 \\

\sigma= 2"

a.


"x= \mu +Z \times \sigma \\

Z=2 \\

x = 46 + 2 \times 2 = 50"

b.

"P(X<42) =P(Z< \frac{42-46}{2}) \\

= P(Z< -2) \\

= 0.0227 \\

= 2.27 \; \%"

c.

"P(46<X<48) = P(X<48) -P(X<46) \\

= P(Z< \frac{48-46}{2}) -P(Z< \frac{46-46}{2}) \\

= P(Z< 1) -P(Z< 0) \\

= 0.8413 -0.5 \\

= 0.3413"

d. Incomplete problem.